# Quantitative surgery and total mean curvature

### v1

@@ 5. Surgery bending in arbitrary backgrounds
In this section, we will construct the general surgery bending. To this end, let $S$ be a compact submanifold of codimension $k$ in an $n$-dimensional Riemannian manifold $(M,g)$. We denote its normal bundle by $\mathcal{N}_{S}$. Recall the constant $R\in(0,1]$ from the previous section, which we may choose smaller if necessary, so that the normal exponential map of $\mathcal{N}_{S}$ along $S$ is a diffeomorphism on the sub-bundle of vectors with length less than $R$. By abuse of notation, we call this sub-bundle $\mathcal{N}_{S}$ as well, and we denote by $D(S)$ the tubular neighbourhood of $S$ given by the image of $\mathcal{N}_{S}$ under the normal exponential map.
In order to define the submanifold $\Sigma_{\varrho}\subset M\times[0,1]$ we first choose Fermi coordinates around $S$, that is, we have an identification $U\times B^{n-k}_{R}(0)\cong D(S)|_{U}=\exp^{\perp}(\mathcal{N}_{S}|U)$, where $U\subseteq S$ is an open set over which $\mathcal{N}_{S}$ is trivial. We define for $\varrho\in(0,R^{2}/4)$
@@
for $\overline{\Sigma}_{\varrho}\subset B^{n-k}_{1}(0)\times[0,1]$ as in 4.1. Note that the transition maps of our local trivialisation are given by families of orthogonal matrices acting on the normal fibers. Thus, by the O(nk)O(n-k)-invariance of $\overline{\Sigma}_{\varrho}$ (cf. Proposition 4.3, (i)), the local definitions can be glued together into a global definition of $\Sigma_{\varrho}\subset M\times[0,1]$. Moreover, since $2\sqrt{\varrho}<R$, the hypersurface $\Sigma_{\varrho}$ can be extended by $(M\setminus D(S))\times\{0\}$ to a smooth hypersurface in $M\times[0,1]$.
The main result of this section is the following theorem, which states the existence of a background metric on $D(S)\times[0,1]$ that extends $g$ on $D(S)\times\{0\}$ such that for $\varrho$ small enough, the scalar curvature of $\Sigma_{\varrho}$ can be uniformly bounded in terms of the scalar curvature of the original metric $g$ and for $\varrho$ even smaller the mean curvature of $\Sigma_{\varrho}$ is positive.
@@
Let $n-k\geq 3$. Then, there exists a constant $0<\varrho_{1}<R^{2}/4$ such that for every $\varrho\in(0,\varrho_{1}]$, every $\mu\in(0,1]$ and for the metric
$g_{\mu}=u^{\frac{4}{\,n-2\,}}(g+dt^{2}),\qquad u(x,t)=u_{\mu}(x,t)=1+t\mu\eta\left(3t-1\right)$
on $D(S)\times[0,1]$ the scalar curvature of the induced metric $g_{\mu,\varrho}$ on $\Sigma_{\varrho}$ satisfies
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq\mathrm{scal}(M,g)-\varrho^{1/4}.$
Furthermore, there exists a constant $\varrho_{0}\in(0,\varrho_{1})$ such that the mean curvature of $\Sigma_{\varrho}$ with respect to $g_{\mu}$ is positive for $\varrho\in(0,\varrho_{0}]$.
@@
    We emphasize that $\varrho_{1}$ is independent of $\mu$, while $\varrho_{0}$ will depend on $\mu$.
@@
    The mean curvature of the slice $M\times\{0\}\subset M\times[0,1]$ with respect to the ambient metric $g_{\mu}$ is given by
    $H(M\times\{0\},g_{\mu})=\frac{2n}{n-2}\mu.$
@@
    If nk=2n-k=2 and $n\geq 3$, then the second statement of Theorem 5.1 still holds true, that is, the mean curvature of $\Sigma_{\varrho}$ with respect to $g_{\mu}$ is still positive. However, the scalar curvature estimate might be incorrect.
Essentially, the content of Theorem 5.1 is that the handle construction of Proposition 4.3 extends to arbitrary Riemannian manifolds and to surgeries of any codimension $n-k\geq 3$. This leads to a number of error terms. Fortunately, they can be grouped into three different categories:
@@
    Error terms of the form $\epsilon|\overline{A}_{\varrho}|$, $\epsilon\ll 1$, where we note that the second fundamental form $|A_{\varrho}|$ blows up as $\varrho\to 0$. Using the pinching estimates $H(\overline{\Sigma}_{\varrho})>\tfrac{1}{2}|A(\overline{\Sigma}_{\varrho})|$ and $\mathrm{scal}(\overline{\Sigma}_{\varrho})>\tfrac{1}{3}|A(\overline{\Sigma}_{\varrho})|^{2}$ we can ensure that they can be absorbed.
@@
There exists a constant $\varepsilon>0$ such that for every $\varrho\leq\varepsilon$ and for the metric $g_{\varrho}$ on $\Sigma_{\varrho}$ induced by $g+\mathrm{d}t^{2}$, the following inequality holds:
(5.1) $\mathrm{scal}(\Sigma_{\varrho},{}$ $g_{\varrho})-\mathrm{scal}(M,g)-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\gtrsim-(\varrho+\|\overline{\nu}_{\varrho}^{\intercal}\|)\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}$
Here, $\|\overline{\nu}_{\varrho}^{\intercal}\|$ denotes the norm of the tangential part of the normal vector of the hypersurface $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{n-k+1}$ constructed in 4.1 with respect to the Euclidean metric (see also (4.1)), and $\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$ denotes the scalar curvature of this hypersurface with respect to the metric $g_{\varrho,\delta}$ induced by the Euclidean ambient metric $\delta+\mathrm{d}t^{2}$. Furthermore, we have the following inequality concerning the mean curvature
(5.2) $|H(\Sigma_{\varrho},g+\mathrm{d}t^{2})-H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\lesssim\|\overline{\nu}_{\varrho}^{\intercal}\|+\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|_{\delta+\mathrm{d}t^{2}}),$
and we can compare the $\partial_{t}$-part of the normal vector $\nu_{\varrho}$ of $\Sigma_{\varrho}$ with respect to $g+\mathrm{d}t^{2}$ and the normal vector $\overline{\nu}_{\varrho}$ of $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{n-k+1}$ taken with respect to the Euclidean metric $\delta+\mathrm{d}t^{2}$ as follows:
@@
Let $u$, $g_{\mu}$ and $g_{\varrho,\mu}$ be as defined in Theorem 5.1. Then we have the following equalities comparing the scalar and mean curvature before and after conformal transformation by $u^{4/(n-2)}$:
@@
(5.5) $\begin{split}H(\Sigma_{\varrho},g_{\mu})=u^{-\frac{2}{n-2}}\left(H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u\right)\end{split}$
@@
We define $\varrho_{1}$ as follows:
$\varrho_{1}\coloneqq{}\frac{1}{2}\min\biggl\{$ $\varepsilon,\,\frac{1}{10},\,\frac{1}{36C},\,\frac{1}{12\cdot(33|\mathrm{scal}(M,g)|+360(n-1))},\,$
$\frac{1}{(9C\overline{C})^{2}},\,\frac{1}{2C\overline{C}(14n-5)},\frac{1}{(20|\mathrm{scal}(M,g)|^{2})},\,\frac{1}{8C\overline{C}+4(n-1)+2C}\biggr\}$
where $C$ and $\varepsilon$ are the constants from Lemma 5.3 and $\overline{C}=18(n-k)+432$ is the constant from Proposition 4.3.
In order to estimate the scalar curvature and mean curvature of $g_{\varrho,\mu}$, we first note that at points with radial coordinate $r\geq 2\sqrt{\varrho}$ we have $\Sigma_{\varrho}=M$ and $g_{\varrho}=g$ So, the scalar curvature estimate is trivially true here, whereas the estimate on the mean curvature follows from Lemma 5.4 using $\partial_{t}u=\mu$, $u=1=\langle\nu_{\varrho},\partial_{t}\rangle$ and $H(M\times\{0\},g+\mathrm{d}t^{2})=0$. Hence, it suffices to consider $r\leq 2\sqrt{\varrho}$. We will distinguish $3$ cases: $r\in[\varrho^{4},\varrho^{2}]$, $r\in[\varrho^{2},\sqrt{\varrho}]$ and $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$. In every case we will first estimate the scalar and mean curvature with respect to the ambient metric $g+\mathrm{d}t^{2}$ using Lemma 5.3 before investigating the effect of conformal change on scalar curvature employing Lemma 5.4. At the very end of the proof, we will prove positivity of the mean curvature after conformal change after passing to a possibly smaller $\varrho_{0}$, depending on $\mu$.
@@
$\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\geq\frac{1}{2}\varrho^{-3}\quad\text{and}\quad\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\geq\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2},$
which also yields $\mathrm{scal}(\overline{\Sigma}_{\varrho})\geq\frac{1}{6}\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|$. Furthermore, we can estimate $\|\overline{\nu}_{\varrho}^{\intercal}\|\leq 1$ and $\varrho\leq 1$, and we thus obtain from (5.1) from Lemma 5.3:
(5.6) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)\\ &\qquad={}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-C(\varrho+\|\overline{\nu}_{\varrho}^{\intercal}\|)\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}\\ &\qquad={}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}\\ &\qquad\geq{}\frac{1}{3}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C+\frac{1}{3}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-4C|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\\ &\qquad\qquad+\frac{1}{3}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C\varrho|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\\ &\qquad\geq{}\varrho^{-3}\underbrace{(\frac{1}{6}-2C\varrho^{3})}_{\geq\frac{1}{6}-2C\varrho\geq\frac{1}{12}}+\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\underbrace{\left(\frac{1}{18}-4C\varrho^{3/2}\right)}_{\geq(\frac{1}{6}-4C\varrho)\geq 0}\\ &\qquad\qquad+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\underbrace{\left(\frac{1}{9}-2C\varrho^{2}\right)}_{\geq\frac{1}{9}-2C\varrho\geq 0}\\ &\qquad\geq{}\frac{1}{12}\varrho^{-3}.\end{split}$
For the mean curvature, we have
$H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})\geq\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\quad\text{and}\quad H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})\geq\frac{1}{2}\varrho^{-3/2},$
by Proposition 4.3. Therefore, (5.2) becomes:
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\geq{}$ $H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})-2C-C\varrho\cdot|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|_{\delta+\mathrm{d}t^{2}}$
@@
$\geq{}$ $\underbrace{\frac{1}{4}\varrho^{-3/2}-2C}_{\geq\frac{1}{8}\varrho^{-3/2}}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\underbrace{\left(\frac{1}{4}-C\varrho\right)}_{\geq 0}\geq\frac{1}{8}\varrho^{-3/2}.$

### v2

@@ 6. Surgery bending in arbitrary backgrounds
In this section, we will construct the general surgery bending. To this end, let $\mathcal{S}$ be a compact submanifold of codimension $k$ in an $n$-dimensional Riemannian manifold $(M,g_{M})$. We denote its normal bundle by $\mathcal{N}_{\mathcal{S}}$. Recall the constant $R\in(0,\tfrac{2}{\sqrt{10}}]$ from the previous section, which we may choose smaller if necessary, so that the normal exponential map of $\mathcal{N}_{\mathcal{S}}$ along $\mathcal{S}$ is a diffeomorphism on the sub-bundle of vectors with length less than or equal to $R$ to the tubular neighborhood $B_{R}(\mathcal{S})$ of radius $R$ around $\mathcal{S}$. With this identification, we consider $B_{R}(\mathcal{S})$ as a fibre bundle over $\mathcal{S}$ with fibre $B_{r}^{k}(0)$.
We will construct a smooth submanifold $\Sigma_{\varrho}\subset M\times[0,1]$ describing a surgery bending in $\mathcal{N}_{\mathcal{S}}$ which for $M=\mathbb{R}^{k}$ with the Euclidean metric and $\mathcal{S}$ a point recovers 5.1. First, we choose Fermi coordinates around $\mathcal{S}$, that is, we have an identification $U\times B^{k}_{R}(0)\cong B_{R}(\mathcal{S})|_{U}$, where $U\subseteq\mathcal{S}$ is a trivialising open subset for $\mathcal{N}_{\mathcal{S}}$. We define for $\varrho\in(0,R^{2}/4)$
@@
for the smooth submanifold $\overline{\Sigma}_{\varrho}\subset B^{k}_{R}(0)\times[0,1]$ as in 5.1. Since $\mathcal{N}_{\mathcal{S}}$ has structure group $\mathrm{O}(k)$, using the $\mathrm{O}(k)$-invariance of $\overline{\Sigma}_{\varrho}$ (cf. Proposition 5.3, (i)), the local definitions can be glued together into a global definition of $\Sigma_{\varrho}\subset B_{R}(\mathcal{S})\times[0,1]$. Moreover, since $2\sqrt{\varrho}<R$, the hypersurface $\Sigma_{\varrho}$ can be extended by $(M\setminus B_{R}(\mathcal{S}))\times\{0\}$ to a smooth hypersurface $\Sigma_{\varrho}\subset M\times[0,1]$.
The main result of this section is the following theorem, which states the existence of a background metric on $M\times[0,1]$ that extends $g_{M}$ on $M\times\{0\}$ such that for $\varrho$ small enough, the scalar curvature of $\Sigma_{\varrho}$ can be uniformly bounded in terms of the scalar curvature of the original metric $g_{M}$ and, for $\varrho$ even smaller the mean curvature of $\Sigma_{\varrho}$ is positive.
@@
Let $k\geq 3$ and $\alpha\in(0,\tfrac{1}{3}]$. Then, there exists a constant $0<\varrho_{1}<\min(R^{2}/4,\tfrac{\alpha}{3})$ such that for all $\kappa,\varrho\in(0,\varrho_{1}]$ and for the metric
$g_{\kappa}=u^{\frac{4}{\,n-2\,}}(g_{M}+dt^{2}),\qquad u(x,t)=u_{\kappa}(x,t)=1+t\tilde{\kappa}\eta\left(\alpha^{-1}t-1\right)$
on $M\times[0,1]$, where $\tilde{\kappa}\coloneqq\tfrac{n-2}{2n}\kappa$, the scalar curvature of the induced metric $g_{\kappa,\varrho}$ on $\Sigma_{\varrho}$ satisfies
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho_{1}^{1/4},$
where $\mathrm{pr}_{M}\colon M\times[0,1]\to M$ is the projection. Furthermore, there exists a constant $\varrho_{0}\in(0,\varrho_{1})$ such that the mean curvature of $\Sigma_{\varrho}$ with respect to $g_{\kappa}$ satisfies
$H(\Sigma_{\varrho},g_{\kappa})>(1-\alpha)\kappa-\varrho^{1/4}$
for all $0<\varrho\leq\varrho_{0}$.
@@
    We emphasize that $\varrho_{1}$ is independent of $\kappa$, while $\varrho_{0}$ will depend on $\kappa$.
@@
    The mean curvature of the slice $M\times\{0\}\subset M\times[0,1]$ with respect to the ambient metric $g_{\kappa}$ and the exterior normal is given by
    $H(M\times\{0\},g_{\kappa})=-\frac{2n}{n-2}\tilde{\kappa}=-\kappa.$
@@
    If $k=2$ and $n\geq 3$, then the second statement of Theorem 6.1 still holds true, that is, the mean curvature of $\Sigma_{\varrho}$ with respect to $g_{\kappa}$ is still positive. However, the scalar curvature estimate might be incorrect.
Remark 6.3.
Essentially, the content of Theorem 6.1 is that the handle construction of Proposition 5.3 extends to arbitrary Riemannian manifolds and to surgeries of any codimension $k\geq 3$. This leads to a number of error terms. They can be grouped into three different categories:
@@
    Error terms of the form $\varepsilon|\overline{A}_{\varrho}|$, $\varepsilon\ll 1$, where we note that the second fundamental form $|A_{\varrho}|$ blows up as $\varrho\to 0$. Using the pinching estimates $H(\overline{\Sigma}_{\varrho})>\tfrac{1}{2}|A(\overline{\Sigma}_{\varrho})|$ and $\mathrm{scal}(\overline{\Sigma}_{\varrho})>\tfrac{1}{3}|A(\overline{\Sigma}_{\varrho})|^{2}$ we can ensure that they can be absorbed.
@@
There exists a constant $\varepsilon>0$ such that for every $0<\varrho\leq\varepsilon$ and for the metric $g_{\varrho}$ on $\Sigma_{\varrho}$ induced by $g+\mathrm{d}t^{2}$, the following inequality holds:
(6.1) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\\ &\qquad\qquad\gtrsim-\varrho\cdot\bigl(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\bigr)^{2}-\|\overline{\nu}_{\varrho}^{\intercal}\|\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\end{split}$
Here, $\|\overline{\nu}_{\varrho}^{\intercal}\|$ denotes the norm of the tangential part of the normal vector of the hypersurface $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{k+1}$ constructed in 5.1 with respect to the Euclidean metric (see also (5.1)), and $\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$ denotes the scalar curvature of this hypersurface with respect to the metric $g_{\varrho,\delta}$ induced by the Euclidean ambient metric $\delta+\mathrm{d}t^{2}$, composed with the projection onto the fiber of the tubular neighbourhood $D(\mathcal{S})$. Furthermore, we have the following inequality concerning the mean curvature
(6.2) $|H(\Sigma_{\varrho},g+\mathrm{d}t^{2})-H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\lesssim\|\overline{\nu}_{\varrho}^{\intercal}\|+\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|),$
and we can compare the $\partial_{t}$-part of the normal vector $\nu_{\varrho}$ of $\Sigma_{\varrho}$ with respect to $g+\mathrm{d}t^{2}$ and the normal vector $\overline{\nu}_{\varrho}$ of $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{k+1}$ taken with respect to the Euclidean metric $\delta+\mathrm{d}t^{2}$ as follows:
@@
Let $u$, $g_{\kappa}$ and $g_{\varrho,\kappa}$ be as defined in Theorem 6.1. Then we have the following equalities comparing the scalar and mean curvature before and after conformal transformation by $u^{4/(n-2)}$:
@@
(6.5) $\begin{split}H(\Sigma_{\varrho},g_{\kappa})=u^{-\frac{2}{n-2}}\left(H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u\right)\end{split}$
@@
In order to estimate the scalar curvature and mean curvature of $g_{\varrho,\kappa}$, we first note that at points with radial coordinate $r\geq 2\sqrt{\varrho}$ we have $\Sigma_{\varrho}=M$ and $g_{\varrho}=g$. So, the scalar curvature estimate is trivially true here, whereas the estimate on the mean curvature follows from Lemma 6.5 using $\partial_{t}u=\tilde{\kappa}$, $u=1=\langle\nu_{\varrho},\partial_{t}\rangle$ and $H(M\times\{0\},g+\mathrm{d}t^{2})=0$. Hence, it suffices to consider $r\leq 2\sqrt{\varrho}$. We will distinguish $3$ cases: $r\in[\varrho^{4},\varrho^{2}]$, $r\in[\varrho^{2},\sqrt{\varrho}]$ and $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$. In every case we will first estimate the scalar and mean curvature with respect to the ambient metric $g+\mathrm{d}t^{2}$ using Lemma 6.4 before investigating the effect of conformal change on scalar curvature employing Lemma 6.5. At the very end of the proof, we will prove positivity of the mean curvature after conformal change after passing to a possibly smaller $\varrho_{0}$, depending on $\kappa$.
@@
$\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})>\frac{1}{2}\varrho^{-3}\quad\text{and}\quad\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})>\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2},$
where $g_{\varrho,\delta}$ is the metric on $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{k+1}$ induced by the ambient Euclidean metric $\delta$. This also yields $\mathrm{scal}(\overline{\Sigma}_{\varrho})\geq\frac{1}{6}\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|$. Hence,
@@
$\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})>\frac{1}{20}\left(\varrho^{-3}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}+\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\right).$
Furthermore, we obtain from (6.1) from Lemma 6.4 that there exist constants $C_{0}>0$, independent of $\kappa$, such that:
(6.6) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}\\ &\quad>{}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-C_{0}\Bigl(\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}+\|\overline{\nu}_{\varrho}^{\intercal}\|(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\Bigr),\end{split}$
where $\mathrm{pr}_{M}\colon\Sigma_{\varrho}\subset M\times[0,1]\to M$ denotes the projection. We will omit $\circ\mathrm{pr}_{M}$ for the remainder of this proof to ease the notation. Since $\|\overline{\nu}_{\varrho}^{\intercal}\|\leq 1$, and we thus obtain
$\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})\\ &\qquad>\frac{1}{20}\left(\varrho^{-3}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}+\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\right)\\ &\qquad\qquad-C_{0}\Bigl(\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}+(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\Bigr)\\ &\qquad\geq\varrho^{-3}\left(\frac{1}{20}-C_{0}\varrho^{3}\right)+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\left(\frac{1}{20}\varrho^{-3/2}-C_{0}\right)\\ &\qquad\qquad+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\left(\frac{1}{20}-C_{0}\varrho\right)\end{split}$
Hence, for $c_{0}\coloneqq\tfrac{1}{40}$ and $\varrho_{1}\leq\tfrac{1}{40C_{0}}$, we have
(6.7) $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})>\mathrm{scal}(M,g_{M})+c_{0}\cdot\varrho^{-3}.$
For the mean curvature, we first recall that by (6.2) from Lemma 6.4 there exists a constant $C_{1}$ such that
(6.8) $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})-H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})\geq-C_{1}\left(\|\overline{\nu}_{\varrho}^{\intercal}\|+\varrho(1+|A(\overline{\Sigma},\delta+\mathrm{d}t^{2})|)\right).$
Furthermore, we have
$H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})>\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\quad\text{and}\quad H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})>\frac{1}{2}\varrho^{-3/2},$
by Proposition 5.3. Therefore, $H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})>\tfrac{1}{4}(|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|+\varrho^{-3/2})$ and (6.8) becomes:
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>{}$ $\frac{1}{4}\left(\varrho^{-3/2}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\right)$
$-C_{1}\left(1+\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\right)$
$={}$ $\varrho^{-3/2}\left(\frac{1}{4}-C_{1}(\varrho^{3/2}+\varrho^{5/2})\right)+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\left(\frac{1}{4}-C_{1}\varrho\right)$
Hence, for $c_{1}\coloneqq\tfrac{1}{8}$ and $\varrho_{1}\leq\tfrac{1}{16C_{1}}$, we get
(6.9) $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>c_{1}\cdot\varrho^{-3/2}$